OpenStudy (alexislecuona1):

At a softball game the concession stand had 125 hot dogs and sold 120 of them. What percent was sold?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (alexislecuona1):

Not too sure

OpenStudy (skullpatrol):

$$\Huge \dfrac{x}{100}=\dfrac{a}{b}$$

OpenStudy (skullpatrol):

Does that^ formula look familiar?

OpenStudy (skullpatrol):

$$\Huge \dfrac{x}{100}=\dfrac{a}{b}$$ Let b = 125 and a = 120

OpenStudy (skullpatrol):

$$\Huge \dfrac{x}{100}= x•\dfrac{1}{100}=x\text{%}=\dfrac{120}{125}$$

OpenStudy (skullpatrol):

$$x•\dfrac{1}{100}=\dfrac{120}{125}$$ $$x•\dfrac{1}{100}=\dfrac{2•2•2•3•5}{5•5•5}$$ $$x•\dfrac{1}{100}=\dfrac{2•2•2•3}{5•5}$$ $$x•\dfrac{1}{100}=\dfrac{24}{25}$$ $$x =?$$

OpenStudy (skullpatrol):

recall: $$\Huge \dfrac{1}{100}=\text{%}$$